OpenStudy (anonymous):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)). So far I have, f(g(x)) = f(-x+3) = 9x(-x + 3) -2 = (-9x^2 + 27x) -2 - what am I doing wrong because I know I messed up somewhere.

hartnn (hartnn):

f(-x+3) = 9x(-x + 3) -2 no! where did the extra 'x' come from ??? f(-x+3) = 9(-x + 3) -2 right??

hartnn (hartnn):

because you just replace 'x' by -x+3

OpenStudy (anonymous):

Oh gosh I feel stupid now. Thanks!

hartnn (hartnn):

don't bother, silly mistakes do happen :) welcome ^_^

hartnn (hartnn):

oh, and since you are new here \[ \begin{array}l\color{red}{\text{W}}\color{orange}{\text{E}}\color{#e6e600}{\text{L}}\color{green}{\text{C}}\color{blue}{\text{O}}\color{purple}{\text{M}}\color{purple}{\text{E}}\color{red}{\text{ }}\color{orange}{\text{t}}\color{#e6e600}{\text{o}}\color{green}{\text{ }}\color{blue}{\text{O}}\color{purple}{\text{P}}\color{purple}{\text{E}}\color{red}{\text{N}}\color{orange}{\text{S}}\color{#e6e600}{\text{T}}\color{green}{\text{U}}\color{blue}{\text{D}}\color{purple}{\text{Y}}\color{purple}{\text{!}}\color{red}{\text{!}}\color{orange}{\text{}}\end{array} \] ask me if you have any doubts browsing this site :)